OpenStudy (anonymous):

Write the function in slope-intercept form. 2x - 5y = -6

OpenStudy (anonymous):

so you want to isolate the Y. first i would subtract 2x to both sides. what do you get? :)

OpenStudy (anonymous):

@LilVeroDez13 you get it so far? :)

OpenStudy (anonymous):

you would get -5y = -6 - 2x.. right?

OpenStudy (anonymous):

yes! but instead just move the -2x in front of the -6. your putting it in slope intercept form, remember ;) now just isolate Y! divide both sides by -5.

OpenStudy (anonymous):

ah yes :) so -5y = -2x - 6 divide by -5 and get y = somethin.. would you divide both -2 and 6 by -5

OpenStudy (anonymous):

yup! you "divide" them both. i say "divide" because you dont have to solve it..make sense?

OpenStudy (anonymous):

you end up with y=2/5x+6/5...is that an option? or is this a problem where you have to show your work?

OpenStudy (anonymous):

the slope you can keep as a fraction because you do rise over run, but the y-intercept is typically a solid number.

OpenStudy (anonymous):

Yes makes complete sense now :) thank you.

OpenStudy (anonymous):

okay great!

OpenStudy (anonymous):

I was jus makin it harder than it really is :)